OpenStudy (anonymous):

Which of the following represents the linear equation 2x = y + 3 in slope−intercept form?

OpenStudy (anonymous):

Which of the following?

OpenStudy (anonymous):

a. 2x + y = 3 b. 2x − y = 3 c. y = 2x − 3 d. y = 2x + 3

OpenStudy (anonymous):

Its C

OpenStudy (anonymous):

thank you:)